# FedRTS: Federated Robust Pruning via Combinatorial Thompson Sampling

### v1

@@ Abstract
Federated Learning (FL) enables collaborative model training across distributed clients without data sharing, but its high computational and communication demands strain resource-constrained devices. While existing methods use dynamic pruning to improve efficiency by periodically adjusting sparse model topologies while maintaining sparsity, these approaches suffer from issues such as greedy adjustments, unstable topologies, and communication inefficiency, resulting in less robust models and suboptimal performance under data heterogeneity and partial client availability. To address these challenges, we propose Federated Robust pruning via combinatorial Thompson Sampling (FedRTS), a novel framework designed to develop robust sparse models. FedRTS enhances robustness and performance through its Thompson Sampling-based Adjustment (TSAdj) mechanism, which uses probabilistic decisions informed by stable, farsighted information instead of deterministic decisions reliant on unstable and myopic information in previous methods. Extensive experiments demonstrate that FedRTS achieves state-of-the-art performance in computer vision and natural language processing tasks while reducing communication costs, particularly excelling in scenarios with heterogeneous data distributions and partial client participation. Our codes are available at: https://github.com/Little0o0/FedRTS.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Illustration of topology adjustment in existing baselines and FedRTS. Left: Existing baselines adjust model topology based on myopic and unstable aggregated weights and gradients via deterministic magnitude pruning and reactivating, resulting in greedy adjustment and high communication overhead. Right: FedRTS introduces Thompson Sampling-based Adjustment (TSAdj) to adjust the topology based on farsighted and stable prior probability distributions, achieving a robust topology and low communication overhead.
@@
Federated learning (FL) (McMahan et al., 2017; Li et al., 2019; Kairouz et al., 2021) enables edge devices with private local data to collaboratively train a model without sharing data. However, substantial computational and communication demands for training deep learning models often exceed the resource limitations of edge devices in cross-device FL scenarios. Although neural network pruning (Han et al., 2015; Molchanov et al., 2019a; Ma et al., 2021) offers a promising solution by removing redundant parameters, traditional pruning methods depend on resource-intensive dense model training, rendering them impractical for privacy-preserved and resource-constrained FL environments.
@@
To address these challenges, recent federated pruning frameworks (Bibikar et al., 2022; Qiu et al., 2022; Tian et al., 2024; Jiang et al., 2022; Huang et al., 2024, 2023; Munir et al., 2021) have adopted dynamic pruning techniques (Evci et al., 2020; Raihan & Aamodt, 2020; Jayakumar et al., 2020) within FL. These frameworks employ a two-loop training process: in the inner loop, model weights are updated through standard FL rounds with fixed model topology; in the outer loop, the server adjusts the model topology by pruning and reactivating parameters (Evci et al., 2020), as illustrated in Fig. 1 (left). This iterative process generates a specialized sparse model without dense model training, significantly reducing computational costs.
@@
Despite these advancements, existing frameworks suffer from three critical challenges in model topology adjustment (Bibikar et al., 2022; Qiu et al., 2022; Tian et al., 2024; Jiang et al., 2022; Huang et al., 2024, 2023; Munir et al., 2021), as illustrated in Fig. 1 (left). 1). Greedy adjustment: Current methods rely on myopic, aggregated information from a small subset of participating clients, ignoring data from the majority of unseen clients and prior knowledge. This leads to greedy adjustments and reduced robustness (Kairouz et al., 2021). 2). Unstable topology: Deterministic adjustments based solely on aggregated information are prone to instability due to heterogeneous data distributions, resulting in unstable model topologies. 3). Communication inefficiency: Transmitting extensive auxiliary data (e.g., full gradient matrices) for topology updates imposes high communication costs. These limitations hinder the ability of current methods to handle client availability, data heterogeneity, and communication costs effectively, ultimately leading to suboptimal performance and inefficient resource utilization.
@@
To address these challenges, we reframe federated pruning as a combinatorial multi-armed bandit (CMAB) (Chen et al., 2013) problem and identify the above issues located in myopic observation and deterministic decision. Building on this insight, we propose Federated Rubust pruning via combinatorial Thompson Sampling (FedRTS), a novel framework designed to derive robust sparse models for cross-device FL, as illustrated in Fig. 2. FedRTS introduces the Thompson Sampling-based Adjustment (TSAdj) mechanism, depicted in Fig. 1 (right), to address the shortcomings of existing methods. First, TSAdj leverages farsighted probability distributions (including prior information from unseen clients) to mitigate the impact of partial client participation. Second, it uses probabilistic decisions based on stable, comprehensive information rather than deterministic adjustments from unstable aggregated information. Third, FedRTS reduces communication overhead by requiring clients to upload only top-gradient indices for adjustments.
@@
We evaluate FedRTS on computer vision and natural language processing datasets, demonstrating its effectiveness in handling client availability and data heterogeneity. FedRTS achieves higher accuracy, better generalization, and lower communication costs compared to state-of-the-art (SOTA) methods. For instance, on the CIFAR-10 dataset with the ResNet-18 model, FedRTS achieves either a 5.1% accuracy improvement or a 33.3% reduction in communication costs compared to SOTA frameworks. To our knowledge, this paper is the first work to analyze federated pruning through a CMAB lens and apply combinatorial Thompson sampling to stabilize topology adjustments in FL.
@@ 2.1 Federated Dynamic Pruning
$\min_{W,m}\sum_{n=1}^{\mathcal{N}}p_{n}f_{n}(W,m,D_{n}),\quad\mathrm{s.t.}\enspace d\leq d^{\prime},$ (1)
[…]
To solve the problem in Eq. 1 under the target density constraint $d^{\prime}$, existing federated pruning frameworks (Bibikar et al., 2022; Qiu et al., 2022; Tian et al., 2024; Jiang et al., 2022; Huang et al., 2024, 2023; Munir et al., 2021) apply dynamic pruning techniques (Evci et al., 2020; Raihan & Aamodt, 2020; Jayakumar et al., 2020) within FL. These frameworks iteratively adjust sparse on-device models during training while maintaining the density level $d$. The process begins with a sparse model initialized via random pruning and follows a two-loop training procedure: In the inner loop, the model topology $m$ remains fixed, and the model weights are updated through traditional FL rounds. After $\Delta T$ inner loops, as illustrated in Fig. 1 (left), the framework enters the outer loop, where clients upload additional data (specifically, the gradients of the pruned weights) to the server. The server then adjusts the model topology by pruning and reactivating (Evci et al., 2020) based on the aggregated weights and gradients.
@@ 2.2 CMAB-Based Problem Formulation
The Combinatorial Multi-Armed Bandit (CMAB) (Chen et al., 2013; Slivkins et al., 2019; Kong et al., 2021) is a sequential decision-making framework where an agent selects combinations of arms (called super arms) with unknown reward distributions. The objective is to maximize cumulative rewards by balancing the exploration of uncertain arms and the exploitation of high-performing ones.
@@
We formulate the topology adjustment as a CMAB problem, where each link $m_{i}$ in the model topology serves as an arm. In each outer loop $t$ for topology adjustment, the server selects an action $S_{t}$, which includes $K$ arms, where $K=d^{\prime}\langle W\rangle$. Selected arms $i\in S_{t}$ are activated ($m_{t,i}=1$), while others $i\notin S_{t}$ are deactivated ($m_{t,i}=0$). After playing action, the new sparse weights $W_{t}=W^{agg}_{t}\odot m_{t}$ will be distributed to the clients for further training, where $W^{agg}_{t}$ is the aggregated weights. The server then observes outcomes $X_{t}=(X_{t,1},\cdots,X_{t,m})$, drawn from distributions with unknown expectation $\mu$. After that, the server obtains the unknown rewards $R_{t}=R(S_{t},X_{t})$. At the next round $t+1$, the previous information $\{X_{\tau}|1\leq\tau\leq t\}$ is the input to the adjustment algorithm to select the action $S_{t+1}$. Following previous work (Kong et al., 2021), we assume that the expected reward of an action $S_{t}$ only depends on $S_{t}$ and and the mean vector $\mu$, i.e., there exists a function $r$ such that $\mathbb{E}[R_{t}]=\mathbb{E}_{X_{t}}[R(S_{t},X_{t})]=r(S_{t},\mu)$.
2.3 Challenges in Previous Methods
@@
From the CMAB aspect, after taking the action $S_{t}$, previous federated pruning methods observe the delay outcomes $X_{t}$ at the next outer loop $t+1$, where $X_{t,i}=1$ if $i$ is the index of the top magnitude of aggregated weights $|W^{agg}_{t+1}|$ or gradients $|G^{agg}_{t+1}|$; otherwise, $X_{t,i}=0$. Treating the outcomes $\{X_{t}\}$ as the feedback, the next action $S_{t+1}$ is selected, i.e., $S_{t+1}=\{i|X_{t,i}=1\}$. Thereby these methods face three significant challenges:
@@
Greedy Adjustment: Existing frameworks only observe the outcomes in the outer loop, ignoring the inner loop. Moreover, they discard all previous information $\{X_{\tau}|1\leq\tau\leq t-1\}$, which is myopic, excluding information from unseen clients and previous data, leading to a greedy topology adjustment that is difficult to exploit from a global view.
@@
Unstable Topology: Previous ignore the expectation of distribution $\mu$ and make a deterministic decision based on unreliable outcomes $X_{t}$. It is hard to maximize the reward expectation $r(S_{t},\mu)$ and may select an arm with a low expectation, leading to an unstable topology.
@@
Communication Inefficiency: Existing frameworks require clients to upload gradients for pruned weights during topology adjustments, making the total communication cost equivalent to that of a dense model, thus lacking communication efficiency. Although some approaches try to mitigate this by uploading only the top gradients (Huang et al., 2023, 2024) or placing adjustments on the client side (Bibikar et al., 2022; Tian et al., 2024), these methods make the adjustment more unstable, leading to suboptimal performance.
@@ 3 Methodology
In this section, we present FedRTS, a novel federated pruning framework designed to develop robust model topologies, as illustrated in Fig. 2. We begin by detailing the TSAdj mechanism, which serves as the core component for searching robust sparse topologies. Next, we provide an overview of the FedRTS algorithm that integrates TSAdj. Finally, we provide the theoretical regret upper bound of TSAdj.
3.1 Thompson Sampling-based Adjustment
We find that the core source of the aforementioned challenges in existing federated dynamic pruning frameworks lies in deterministic decision strategies and myopic observations. Therefore, we propose the Thompson Sampling based Adjustment (TSAdj) mechanism, which includes probabilistic decision strategy and farsighted observations. Moreover, TSAdj effectively reduces the communication overhead.
Compared to directly using myopic outcomes to make deterministic decisions, our proposed TSAdj maintains individual prior probability distributions $P$ for each link in the model topology $m$ during training. The higher feedback a link $i$ has gained from previous observations, the closer the expectation of its probability $E[P_{i}]$ approaches 1, indicating a higher likelihood of being selected. In the outer loop for topology adjustment, instead of deterministically selection based on outcomes $X_{t}$, the server samples random variables $\xi\sim P$ and selects an action $S_{t}$ on the topology as follows:
$S_{t}=\{i|\xi_{i}\in\text{Top}(\xi,K)\},$ (2)
where $\text{Top}(\xi,K)$ denotes the set of top $K$ elements of $\xi$. The action $S_{t}$ is derived from the expectation of the prior distribution, $E[P]$, which provides a more stable topology compared to existing approaches.
Unlike existing approaches that develop the outcomes using immediate and myopic data in the outer loop, TSAdj derives much more comprehensive outcomes $X_{t}$ from both the inner and outer loops. After performing action $S_{t}$, the outcome $X_{t}$ is observed in the end of loop $t$, and is formulated as a combination of individual outcomes $X^{n}_{t}$ and aggregated outcomes $X^{agg}_{t}$:
$X_{t}=(1-\gamma)X^{agg}_{t}+\gamma\sum_{n\in C_{t}}p_{n}X^{n}_{t},$ (3)
@@ 3.1 Thompson Sampling-based Adjustment
We consider the semi-outcomes for $X^{agg}_{t}$ and $X^{n}_{t}$ in the inner loop, i.e., only the arms $i\in S_{t}$ have the outcomes. We define a constant $\kappa<K$, which denotes the number of core links (arms) in the model topology, and during the adjustment, we expect that $K-\kappa$ less important links are pruned, while $K-\kappa$ candidate links are reactivated. Therefore, when the loop $t+1$ is the inner loop, we have
$X^{agg}_{t,i}=h(i,|W^{agg}_{t+1}|,\kappa),\quad i\in S_{t},$ (4)
$X^{n}_{t,i}=h(i,|W^{n}_{t+1}|,\kappa),\quad i\in S_{t},$ (5)
where $W^{agg}_{t+1}$ and $W^{n}_{t+1}$ represent the aggregated weights and client $n$’s weights, respectively, and $h(\cdot)$ is a discriminative function defined as $h(i,x,\kappa)=\mathbf{1}_{i\in\text{Top}(x,\kappa)}$. In the inner loop, for arms $i\notin S_{t}$, both the aggregated outcomes $X^{agg}_{t,i}$ and the individual outcomes $X^{n}_{t,i}$ are set to $None$. In the outer loop, we consider the full outcomes of $X^{agg}_{t}$ and $X^{n}_{t}$ as the gradients for inactivated weights are available. For the arms $i\notin S_{t}$, the individual outcomes $X^{n}_{t,i}$ are defined as
$X^{n}_{t,i}=h(i,|G^{n}_{t+1}|,K-\kappa),\quad i\notin S_{t},$ (6)
where the $G^{n}_{t+1}$ denotes the gradients of the weights $W^{n}_{t+1}$. Here, $K-\kappa$ is used to treat some inactivated arms as candidates for reactivation. We set the aggregated outcomes $X^{agg}_{t,i}=0.5$ for $i\notin S_{t}$ in the outer loop, without using the aggregated gradients $G^{agg}_{t+1}$ due to their high variance. Notably, the clients only upload the Top $K-\kappa$ indices of gradients magnitude $|G^{n}_{t+1}|$ to achieve 6, rather than the full gradients, significantly reducing the communication overhead.
To construct a posterior probability distribution, TSAdj updates the prior distribution $P$ based on outcomes $X_{t}$. Following the Thompson Sampling framework, each $P_{i}$ is modeled as a Beta distribution with factors $\alpha_{i},\beta_{i}$, representing the likelihood of selecting an arm into $S_{t}$: a higher $\alpha_{i}$ shifts $\mathbb{E}[P_{i}]$ towards 1, whereas a higher $\beta_{i}$ shifts it towards 0. The $(\alpha_{i},\beta_{i})$ are updated via Bayes’ rule, leveraging its conjugacy properties and using the outcomes $X_{t,i}$ as follows:
@@
  Input: Uploaded sparse weights $\{W^{n}_{t+1}\}_{n\in C_{t}}$ and gradients $\{G_{t+1}^{n}\}_{n\in C_{t}}$, global distribution factors $(\alpha,\beta)$
  Output: Updated weights $W_{t+1}$ and topology $m_{t+1}$
@@
  $X^{agg}_{t}\leftarrow$ calculate by Eq.4 based on $W^{agg}_{t+1}$
  $X^{n}_{t}\leftarrow$ calculate by Eq.5,6 based on $W_{t+1}^{n}$ and $G_{t+1}^{n}$
  $X_{t}\leftarrow(1-\gamma)X^{agg}_{t}+\gamma\sum_{n\in C_{t+1}}p_{n}X^{n}_{t}$
  $\xi\leftarrow$ sample from $Beta(\alpha,\beta)$
  $S_{t}\leftarrow\{i|\xi_{i}\in\text{Top}(\xi,K)\}$
  $m_{t+1}\leftarrow$ adjust $m_{t}$ based on action $S_{t}$
@@
In summary, TSAdj leverages the prior probability distributions $P$ to probabilistically guide topology adjustments and uses comprehensive outcomes $X_{t}$ to update the distributions, which mitigates instability caused by myopic adjustments, enabling smoother convergence and developing a more reliable topology. Moreover, TSAdj requires the clients to upload only the most critical gradient information, thereby reducing the communication overhead.
3.2 FedRTS
FedRTS is a novel federated learning framework designed to address the limitations of existing dynamic sparse training methods by integrating the Thompson Sampling-based Adjustment (TSAdj) mechanism, as illustrated in Fig. 2. The framework begins with an initialized model weight and sparse topology sampled from Beta distributions $P$. FedRTS employs a two-loop training process to iteratively update model weights and refine the sparse topology.
In the inner loop, the model topology remains fixed while the server updates the model weights and distribution $P$ based on semi-outcomes. In the outer loop, the server collects the full outcomes by requiring clients to upload the top-$\kappa$ gradient indices of inactivated weights. Using this information, the server applies TSAdj to make robust topology adjustments, and the updated topology is then distributed to clients. FedRTS continues this iterative process until model convergence. The details of FedRTS are shown in Algo. 28 and Sec. B.2.

### v3

@@ Abstract
Federated Learning (FL) enables collaborative model training across distributed clients without data sharing, but its high computational and communication demands strain resource-constrained devices. While existing methods use dynamic pruning to improve efficiency by periodically adjusting sparse model topologies while maintaining sparsity, these approaches suffer from issues such as greedy adjustments, unstable topologies, and communication inefficiency, resulting in less robust models and suboptimal performance under data heterogeneity and partial client availability. To address these challenges, we propose Federated Robust pruning via combinatorial Thompson Sampling (FedRTS), a novel framework designed to develop robust sparse models. FedRTS enhances robustness and performance through its Thompson Sampling-based Adjustment (TSAdj) mechanism, which uses probabilistic decisions informed by stable and farsighted information, instead of deterministic decisions reliant on unstable and myopic information in previous methods. Extensive experiments demonstrate that FedRTS achieves state-of-the-art performance in computer vision and natural language processing tasks while reducing communication costs, particularly excelling in scenarios with heterogeneous data distributions and partial client participation. Our codes are available at: https://github.com/Little0o0/FedRTS.
@@ 1 Introduction
Figure 1: Illustration of topology adjustment in existing baselines  [^2] Baseline’s workflow: https://github.com/Little0o0/FedRTS/tree/main/image/workflow.mp4and FedRTS. Left: Existing baselines adjust model topology based on myopic and unstable aggregated weights and gradients via deterministic magnitude pruning and reactivating, resulting in greedy adjustment and high communication overhead. Right: FedRTS introduces Thompson Sampling-based Adjustment (TSAdj) to adjust the topology based on farsighted and stable probability distributions, achieving a robust topology and low communication overhead.
@@
Federated learning (FL) [34, 30, 25, 64, 65] enables edge devices with private local data to collaboratively train a model without sharing data. However, substantial computational and communication demands for training deep learning models often exceed the resource limitations of edge devices in cross-device FL scenarios. Although neural network pruning [12, 40, 33] offers a promising solution by removing redundant parameters, traditional pruning methods depend on resource-intensive dense model training, rendering them impractical for privacy-preserved and resource-constrained FL environments.
@@
To address these challenges, recent federated pruning frameworks [1, 47, 57, 24, 19, 18, 42] have adopted dynamic sparse training techniques [11, 48, 22] within FL. These frameworks employ a two-loop training process: in the inner loop, model weights are updated through standard FL rounds with fixed model topology; in the outer loop, the server adjusts the model topology by pruning and reactivating parameters [11], as illustrated in Fig. 2 (left). This iterative process generates a specialized sparse model without dense model training, significantly reducing computational costs.
@@
Despite these advancements, existing frameworks [1, 47, 57, 24, 19, 18, 42] suffer from three critical challenges in model topology adjustment, as illustrated in Fig. 2 (left). (1) Greedy adjustment: Current methods rely on myopic, aggregated information from a small subset of participating clients, ignoring data from the majority of unseen clients and prior knowledge. This leads to greedy adjustments and reduced robustness [25]. (2) Unstable topology: Deterministic adjustments based solely on aggregated information are prone to instability due to heterogeneous data distributions, resulting in unstable model topologies. (3) Communication inefficiency: Transmitting extensive auxiliary data (e.g., full-size gradients) for topology updates imposes high communication costs. These limitations hinder the ability of current methods to handle client availability, data heterogeneity, and communication costs effectively, ultimately leading to suboptimal performance and inefficient resource utilization.
@@
To address these challenges, we reframe federated pruning as a combinatorial multi-armed bandit (CMAB) [5] problem and identify the above issues located in myopic observation and deterministic decision. Building on this insight, we propose Federated Rubust pruning via combinatorial Thompson Sampling (FedRTS), a novel framework designed to derive robust sparse models for cross-device FL, as illustrated in Fig. 2. FedRTS introduces the Thompson Sampling-based Adjustment (TSAdj) mechanism, depicted in Fig. 2 (right), to address the shortcomings of existing methods. First, TSAdj leverages farsighted probability distributions (including prior information from unseen clients) to mitigate the impact of partial client participation. Second, it uses probabilistic decisions based on stable, comprehensive information for topology adjustment. Third, FedRTS reduces communication overhead by requiring clients to upload only top-gradient indices instead of dense gradients.
@@
We evaluate FedRTS on computer vision and natural language processing datasets, demonstrating its effectiveness in handling client availability and data heterogeneity. FedRTS achieves higher accuracy, better generalization, and lower communication costs compared to state-of-the-art (SOTA) methods. For instance, on the CIFAR-10 dataset with the ResNet18 model, FedRTS achieves either a 5.1% accuracy improvement or a 33.3% reduction in communication costs compared to SOTA frameworks. To our knowledge, this paper is the first work to analyze federated pruning through a CMAB lens and apply combinatorial Thompson sampling to stabilize topology adjustments in FL.
@@ 2.1 Federated Dynamic Pruning
$\min_{W,m}\sum_{n=1}^{\mathcal{N}}p_{n}f_{n}(W,m,D_{n}),\quad\mathrm{s.t.}\kern 5.0ptd\leq d^{\prime},$ (1)
where $W$ represents the model weights, $m\in\{0,1\}^{\langle W\rangle}$ denotes the sparse model topology (also called mask), $f_{n}(\cdot)$ is the objective function depending on the task of the $n$-th for client $n$ (e.g., cross-entropy loss for the image classification task), $d=\frac{\mathrm{sum}(m)}{\langle m\rangle}$ is the density of the topology $m$, and $p_{n}$ is set as the proportion of the data size of client $n$, typically set as $p_{n}=\langle D_{n}\rangle/\sum^{\mathcal{N}}_{i=1}\langle D_{i}\rangle$.  [^3] In this paper, $\langle\cdot\rangle$ is the cardinal number of the set.
To solve the problem in Eq. 1 under the target density constraint $d^{\prime}$, existing federated pruning frameworks [1, 47, 57, 24, 19, 18, 42] apply dynamic sparse training techniques [11, 48, 22]. These frameworks iteratively adjust sparse on-device models during training while maintaining the density level $d\leq d^{\prime}$. The process begins with a sparse model initialized via random pruning and follows a two-loop training procedure: In the inner loop, the model topology $m$ remains fixed, and the model weights $W$ are updated through traditional FL rounds. After $\Delta T$ inner loops, as illustrated in Fig. 2 (left), the framework enters the outer loop, where clients upload additional data (specifically, the gradients of the pruned weights) to the server. The server then adjusts the model topology by pruning and reactivating [11] based on the aggregated weights and gradients.
@@ 2.2 CMAB-Based Problem Formulation
The Combinatorial Multi-Armed Bandit (CMAB) [5, 55, 26] is a sequential decision-making framework where an agent selects combinations of arms (called super arms) with unknown reward distributions. The objective is to maximize cumulative rewards by balancing the exploration of uncertain arms and the exploitation of high-performing ones.
@@
We formulate the topology adjustment as a CMAB problem, where each link $m_{i}$ in the model topology serves as an arm. In each outer loop $t$, the server selects an action $S_{t}$, which includes $K$ arms, where $K=d\langle m\rangle$. Selected arms $i\in S_{t}$ are activated, while others $i\notin S_{t}$ are deactivated, i.e., $m_{t,i}=\mathbf{1}_{i\in S_{t}}$. After playing action, the new sparse weights $W_{t}=W^{agg}_{t}\odot m_{t}$ will be distributed to the clients for further training. The server then observes outcomes $X_{t}=(X_{t,1},\cdots,X_{t,\langle m\rangle})$ for all arms, drawn from distributions with unknown expectation $\mu$. After that, the server obtains the unknown rewards $R_{t}=R(S_{t},X_{t})$. At the next round $t+1$, the previous information $\{X_{\tau}|1\leq\tau\leq t\}$ is the input to the adjustment algorithm to select the action $S_{t+1}$. Following previous work [26], we assume that the expected reward of an action $S_{t}$ only depends on $S_{t}$ and and the mean vector $\mu$, i.e., there exists a function $r$ such that $\mathbb{E}[R_{t}]=\mathbb{E}_{X_{t}}[R(S_{t},X_{t})]=r(S_{t},\mu)$. Assuming the number of adjustments is $T$, the goal of CMAB is to maximize the cumulative reward, i.e., $\max\sum_{t=1}^{T}\mathbb{E}[R_{t}]=\sum_{t=1}^{T}r(S_{t},\mu)$.
2.3 Challenges in Previous Federated Dynamic Pruning Methods
@@
From the CMAB aspect, we can rigorously analyse the challenges in the previous methods. After taking the action $S_{t}$ to determine topology $m_{t}$, previous methods observe the delay outcomes $X_{t}$ at the next outer loop $t+1$, defining $X_{t,i}=1$ for $i$-th arm with the top importance score; otherwise, $X_{t,i}=0$. And the next action $S_{t+1}$ is selected merely based on current outcomes $X_{t}$, i.e., $S_{t+1}=\{i|X_{t,i}=1\}$. Thereby, they face three significant challenges:
@@
Greedy Adjustment: Existing frameworks only observe the outcomes in the outer loop, ignoring the inner loop. Moreover, they discard all previous outcomes $\{X_{\tau}|1\leq\tau\leq t-1\}$, excluding information from unseen clients and previous data, leading to a myopic and greedy topology adjustment that is difficult to exploit from a global view.
@@
Unstable Topology: Previous methods make a deterministic decision to select an action $S_{t}$ based on unreliable outcomes only, without considering the high variance of outcomes and the expectation of distribution $\mu$. Therefore, the selected action $S_{t}$ is hard to maximize the reward $r(S_{t},\mu)$, leading to an unstable topology.
@@
Communication Inefficiency: To compute the outcomes $X_{t}$ for all arms, existing frameworks use the magnitudes of aggregated weights and gradients as importance scores for active and inactive arms, respectively. As a result, clients must upload gradients for inactive weights, leading to a total communication cost comparable to that of a dense model, thus lacking communication efficiency. Although some approaches try to mitigate this by uploading only the top gradients [18, 19] or placing adjustments on the client side [1, 57], these greedy methods make the adjustment more unstable, leading to suboptimal performance.
@@ 3 Methodology
In this section, we present FedRTS, a novel federated pruning framework designed to develop robust model topologies, as illustrated in Fig. 2. We begin by detailing the TSAdj mechanism, which serves as the core component for searching robust sparse topologies. Next, we provide an overview of FedRTS that integrates TSAdj. Finally, we provide the theoretical regret upper bound of TSAdj.
3.1 Thompson Sampling-based Adjustment (TSAdj)
We find that the core source of the aforementioned challenges in existing methods lies in deterministic decision strategies and myopic observations. Therefore, we propose the Thompson Sampling-based Adjustment (TSAdj) mechanism, which includes a probabilistic decision strategy based on farsighted observations. Moreover, TSAdj effectively reduces communication overheads.
Compared to directly using myopic outcomes $X_{t}$ to make deterministic decisions, our proposed TSAdj maintains individual posterior probability distributions $P$ for each link in the model topology $m$ during training. The higher feedback a link $i$ has gained from previous observations, the closer the expectation of its probability $E[P_{i}]$ approaches 1, indicating a higher likelihood of being selected. In the outer loop for topology adjustment, the server samples random variables $\xi$, where $\xi_{i}\sim P_{i}$, and selects an action $S_{t}$ as follows:
$S_{t}=\{i\in\text{Top}(\xi,K)\},$ (2)
where $\text{Top}(\xi,K)$ are the indices of top-$K$ elements in $\xi$. The action $S_{t}$ is determined by the posterior distributions $P$, rather than deterministically based on outcomes, thereby providing a more stable topology with lower variance [51] compared to existing approaches.
Unlike existing approaches that develop the outcomes using immediate and myopic data in the outer loop, TSAdj derives much more comprehensive outcomes $X_{t}$ from both the inner and outer loops. After performing action $S_{t}$, the outcome $X_{t}$ is observed in the end of loop $t$, and is formulated as a fusion of individual outcomes $X^{n}_{t}$ and aggregated outcomes $X^{agg}_{t}$:
$X_{t}=\gamma X^{agg}_{t}+(1-\gamma)\sum_{n\in C_{t}}p_{n}X^{n}_{t},$ (3)
@@ 3.1 Thompson Sampling-based Adjustment (TSAdj)
We consider the semi-outcomes for $X^{agg}_{t}$ and $X^{n}_{t}$ in the inner loop, i.e., only the arms $i\in S_{t}$ have the outcomes. We define a constant $\kappa<K$, which denotes the number of core links (arms) in the model topology, and during the adjustment, we expect that $K-\kappa$ less important links are pruned, while $K-\kappa$ candidate links are reactivated. Following previous methods, we use the magnitude of weights to get the outcomes for active arms $i\in S_{t}$. Therefore, at the end of inner loop, we have
$X^{agg}_{t,i}=h(i,|W^{agg}_{t+1}|,\kappa),\quad X^{n}_{t,i}=h(i,|W^{n}_{t+1}|,\kappa),\quad i\in S_{t},$ (4)
where $W^{agg}_{t+1}$ and $W^{n}_{t+1}$ represent the aggregated weights and client $n$’s weights, respectively. Here we use $t+1$ instead of $t$ because the local training is finished. $h(\cdot)$ is a discriminative function defined as $h(i,x,\kappa)=\mathbf{1}_{i\in\text{Top}(x,\kappa)}$. In the inner loop, for arms $i\notin S_{t}$, both the aggregated outcomes $X^{agg}_{t,i}$ and the individual outcomes $X^{n}_{t,i}$ are set to $None$.
We consider the full outcomes of $X^{agg}_{t}$ and $X^{n}_{t}$ for outer loop. Beyond the outcomes of active arms in Eq. 4, we estimate the outcomes for inactive arms $i\notin S_{t}$ using the magnitudes of their gradients,
$X^{agg}_{t,i}=0.5,\quad X^{n}_{t,i}=h(i,|G^{n}_{t+1}|,K-\kappa),\quad i\notin S_{t},$ (5)
where the $G^{n}_{t+1}$ denotes the gradients of the weights $W^{n}_{t+1}$. Here, $K-\kappa$ is used to treat some inactivated arms as candidates for reactivation. We do not use the aggregated gradients $G^{agg}_{t+1}$ to compute $X^{agg}_{t,i}$ due to the high variance. Notably, to achieve Eq. 5, the clients only need to upload the Top indices of gradients magnitude, $\mathcal{I}_{t+1}^{n}=\mathrm{Top}(|G^{n}_{t+1}|,K-\kappa)$, rather than the full gradients, significantly reducing the communication overhead.
To construct a posterior probability distribution, TSAdj updates the distribution $P$ based on outcomes $X_{t}$. Following the Thompson Sampling framework, each $P_{i}$ is modeled as a $Beta$ distribution with factors $\alpha_{i},\beta_{i}$, representing the likelihood of selecting an arm into $S_{t}$: a higher $\alpha_{i}$ shifts $\mathbb{E}[P_{i}]$ towards 1, whereas a higher $\beta_{i}$ shifts it towards 0. The $(\alpha_{i},\beta_{i})$ are updated via Bayes’ rule, leveraging its conjugacy properties and using the outcomes $X_{t,i}$ as follows:
@@
Input: Participated clients $C_{t}$, uploaded sparse weights $\{W^{n}_{t+1}\}_{n\in C_{t}}$ and top gradients’ indices $\{\mathcal{I}_{t+1}^{n}\}_{n\in C_{t}}$, trade-off ratio $\gamma$, scaling factor $\lambda$, global distribution factors $(\alpha,\beta)$, number of selected arms $K$, sparse model topology $m_{t}$.
Output: Updated sparse weights $W_{t+1}$ and topology $m_{t+1}$
@@
$X^{agg}_{t},\{X^{n}_{t}\}_{n\in C_{t}}\leftarrow$ calculate with $W^{agg}_{t+1}$, $\{W^{n}_{t+1}\}_{n\in C_{t}}$ and $\{\mathcal{I}_{t+1}^{n}\}_{n\in C_{t}}$ by Eq. 4 and Eq. 5
$X_{t}\leftarrow\gamma X^{agg}_{t}+(1-\gamma)\sum_{n\in C_{t}}p_{n}X^{n}_{t}$
$(\alpha_{i},\beta_{i})\leftarrow(\alpha_{i}+\lambda X_{t,i},\beta_{i}+\lambda(1-X_{t,i}))$,  for $i=1,2,\dots,\langle X_{t}\rangle$ and $X_{t,i}\not=None$
$S_{t+1}\leftarrow\{i|\xi_{i}\in\text{Top}(\xi,K)\}$,  $\xi=\{\xi_{i}|\xi_{i}\sim Beta(\alpha_{i},\beta_{i})\}$  // Select new action
$m_{t+1}\leftarrow\mathbf{1}_{\{i\in S_{t+1}\}}$  // Adjust topology based on selected action
@@
In summary, TSAdj leverages the posterior probability distributions $P$ to probabilistically guide topology adjustments and uses comprehensive outcomes $X_{t}$ to update the distributions, which mitigates instability caused by myopic adjustments, enabling smoother convergence and developing a more reliable topology. Moreover, TSAdj requires the clients to upload only the most critical gradient information, thereby reducing the communication overhead.
3.2 Proposed FedRTS framework
FedRTS is a novel federated pruning framework designed to address the limitations of existing dynamic sparse training methods by integrating the Thompson Sampling-based Adjustment (TSAdj) mechanism, as illustrated in Fig. 2. The framework begins with an initialized model weight and sparse topology sampled from Beta distributions $P$. FedRTS employs a two-loop training process to iteratively update model weights and refine the sparse topology.
In the inner loop, the model topology remains fixed while the server updates the model weights and distribution $P$ based on semi-outcomes. In the outer loop, the server collects the full outcomes by requiring clients to upload the top gradient indices of inactivated weights. Using this information, the server applies TSAdj to make robust topology adjustments, and the updated topology is then distributed to clients. FedRTS continues this iterative process until model convergence. The details of FedRTS are shown in Algorithm 2 and Sec. B.2.